# A Comprehensive Survey of Small Language Models in the Era of Large Language Models: Techniques, Enhancements, Applications, Collaboration with LLMs, and Trustworthiness

### v1

@@ 5.1. Generic-domain SLMs
“The bigger, the better” dominates LLM development, yet these models struggle with on-device processing and energy efficiency, which are vital for privacy, customization, and sustainability. Small language models (SLMs), with fewer parameters, improve computational efficiency in pre-training, fine-tuning, and inference, reducing memory and energy demands—crucial in resource-limited settings. The localized and compact characteristics of SLMs enhance privacy, personalization, and response times, making them suitable for low-power edge devices. Therefore, SLMs are attracting increasing attention, and various models are being developed. Table 8 summarizes current representative generic-domain $40$ SLMs/SLM families. Although all chosen SLMs have similar architectures, they vary in specific training datasets and techniques, with some datasets not being openly available. Taking the latest Llama 3.2 1B models (AI, 2024) in Figure 17 as an example, its parameter size and use of filtered high-quality training data, pruning-based initialization, knowledge distillation pre-training tasks, and training techniques such as Supervised Fine-Tuning (SFT), Rejection Sampling (RS), and Direct Preference Optimization (DPO) distinguish it from others.
@@
From Table 8, we observe several trends in component choices for SLMs: (i) Recent SLMs frequently employ Grouped Query Attention (GQA) in self-attention mechanisms because it can reduce computational complexity. GQA achieves this by sharing query representations across multiple heads while keeping key and value representations separate. This approach aligns with the goals of SLM to enhance efficiency without compromising functionality. (ii) SiLU and SwiGLU are preferred over ReLU as activation functions due to their superior capabilities. SwiGLU’s parameters are learned during training, allowing the model to dynamically adapt to various tasks and datasets, enhancing its flexibility and establishing it as a state-of-the-art choice. Conversely, SiLU is favored in small language models mainly for its computational efficiency, as it introduces no extra parameters. (iii) RMS normalization is more commonly used than traditional layer normalization due to its reduced computational demands. A basic introduction to these options is provided in Section 2.
Apart from component choices, there are notable innovations in architecture for SLMs:
  • The deep and thin architecture (Liu et al., 2024h) highlights that deeper models are more effective than wider ones for improving performance.
  • Embedding sharing (Zhang et al., 2022) is significant because embedding layers make up a large part of the model’s parameters. For instance, with an embedding dimension of 512 and a vocabulary size of 32k, the input and output embedding layers contain 16 million parameters each. These layers account for over 20% of the total parameters in a 125M-parameter model. In sub-billion scale models, there is a return to the concept of sharing embeddings between the input and output layers, using the same weights for both, which makes the model more efficient and compact.
  • Layer sharing (Liu et al., 2024h) is advantageous for small Transformer models, increasing the number of hidden layers without extra storage costs.
@@ 5.1.1. Training Datasets
From Table 8, we can observe a set of widely used training datasets in SLM development. We provide the details below
  • Pile (Gao et al., 2020): It comprises 22 smaller, high-quality diverse corpora from various domains, such as Pile-CC, PubMed Central, ArXiv, GitHub, and FreeLaw, designed to offer a comprehensive foundation for language model training. The dataset contains 207 billion tokens and totals 825 GiB.
@@
  • The Stack (Kocetkov et al., 2022): It contains 6 trillion tokens of source code from over 300 programming languages, useful for developing code-centric AI applications.
@@
From the analysis of these datasets, we can derive several critical insights regarding the development of SLMs: (i) Data Quality: Data quality is paramount in training effective SLMs. Most pre-training routines incorporate sophisticated filtering processes to exclude low-quality content, which may involve removing duplicates or irrelevant samples based on another LLM. For instance, the TinyStories corpus (Eldan and Li, 2023) is specifically curated to be simple and clear, resembling language suited for a three-year-old child, thereby facilitating the training of models that can comprehend and generate uncomplicated narratives effectively. RedPajama-V2 (Computer, 2023) is processed via the CCNet pipeline, including 30B documents with quality signals, and provides IDs for deduplicating documents to create a 20B deduplicated dataset. (ii) Code Data: Source code constitutes a significant component of valuable data for training models, particularly because of its structured nature and logical content. Training on code data enhances a model’s reasoning capabilities and supports its ability to generalize across multiple natural languages, which is crucial for applications requiring robust problem-solving and interpretation skills in diverse coding environments (MA et al., 2024; Aryabumi et al., 2024; Fu and Khot, 2022; Gunasekar et al., 2023)
@@ 5.1.2. Training Algorithms
  • Direct Preference Optimization (DPO) (Rafailov et al., 2024) introduces a novel methodology for optimizing language models based on human preferences, providing a simpler alternative to Reinforcement Learning from Human Feedback (RLHF). Unlike RLHF, DPO does not employ explicit reward modeling or reinforcement learning techniques. Instead, it uses a dynamic weighting mechanism to modify the log probabilities of preferred versus non-preferred responses. This approach helps to avoid model degradation issues commonly encountered with methods that rely solely on probability ratio targets. The DPO loss function is as follows:
@@
    $\mathcal{L}_{DPO}(\pi_{\theta};\pi_{\text{ref}})=-\mathbb{E}_{(x,y_{w},y_{l})\sim D}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}\mid x)}{\pi_{\text{ref}}(y_{w}\mid x)}-\beta\log\frac{\pi_{\theta}(y_{l}\mid x)}{\pi_{\text{ref}}(y_{l}\mid x)}\right)\right]$
@@
    Here, $\pi_{\theta}$ is the policy being optimized, and $\pi_{\text{ref}}$ is the reference policy. The expectation is taken over a dataset $D$, comprising tuples $(x,y_{w},y_{l})$, where $x$ is the input and $y_{w}$, $y_{l}$ are the preferred and non-preferred outputs, respectively. The sigmoid function $\sigma$, and the scaling factor $\beta$ adjust the log-ratio of probabilities from $\pi_{\theta}$ and $\pi_{\text{ref}}$, guiding the model towards producing outputs aligned with human preferences.
@@
  • Reinforcement Learning from Contrast Distillation (RLCD) (Yang et al., 2024b) is designed to calibrate generative SLMs/LLMs towards embodying harmless and beneficial characteristics. The process commences with an unaligned LM and a series of initial prompts. The prompts are systematically modified into two variants $p+$ and $p$-, which are intended to promote and suppress, respectively, specific attributes such as helpfulness and harmlessness. Upon inputting these altered prompts into the LM, outputs $o+$ and $o$- are generated, with $o+$ automatically marked as the preferred response. This automation expedites the training process by eliminating the requirement for subsequent evaluative scoring. The subsequent training adheres to the conventional Reinforcement Learning from Human Feedback (RLHF) framework. The model is trained on this artificially constructed paired preference data to cultivate a preference model. Subsequently, a reward model is extracted from this preference model. Proximal Policy Optimization (PPO) is then employed to refine the alignment of the original LM, ensuring that its outputs more closely reflect human values and expectations. This method effectively fine-tunes the model’s responses, confirming their compliance with established ethical norms and user preferences.
  • Conditioned Reinforcement Learning Fine-Tuning (C-RLFT), proposed by OpenChat (Wang et al., 2024b), aims to improve model performance by effectively incorporating low-quality data during SFT. Data quality plays a crucial role in model outcomes; for instance, OpenChat 6k (Wang et al., 2023b) uses 6k high-quality ShareGPT data annotated by GPT-4 (Achiam et al., 2023), excluding GPT-3.5 annotations, and outperforms Vicuna (Chiang et al., 2023), which uses the full 70k ShareGPT dataset, and Alpaca (Taori et al., 2023) with 50k self-instruct samples. However, OpenChat 6k overlooks the potential of low-quality data, which can contribute knowledge and help the model distinguish between varying quality levels. C-RLFT addresses this by leveraging mixed-quality data with simple, coarse-grained reward labels that differentiate data quality—e.g., expert data with 1 credit and sub-optimal data with 0.1 credits. To mitigate imperfect reward signals, C-RLFT conditions data sources with distinct prompt tokens recognized by the model itself. Additionally, C-RLFT supports reward-free training and eliminates the need for costly human feedback, making it a flexible and efficient fine-tuning method for SLMs/LLMs. A similar strategy, Data Mix (Pfeiffer et al., 2024), involves training on English text in three stages, each with a different mix of data. In each stage, the proportion of noisy web data is gradually reduced in favor of higher-quality data.
@@
  • Explanation Tuning, proposed by Orca (Mukherjee et al., 2023), addresses the limitations of standard instruction-based fine-tuning, which often restricts SLMs to style imitation rather than reasoning. It uses system prompts with instructions to direct GPT-4 to produce detailed explanations or perform step-by-step reasoning. The resulting instructions and the responses are used as a dataset for fine-tuning SLMs to have better ability of reasoning. Specifically, Orac adopts the training dataset sourced from FLAN-v2 (Longpre et al., 2023a), which includes five subsets covering various datasets. Five million questions from FLAN-v2 are initially processed through ChatGPT, and one million of these responses are further refined by GPT-4 to create the training set. Many models such as Orca 1 (Mukherjee et al., 2023), StableBeluga (Mahan et al., [n. d.]) and Dolphin  [^7] https://huggingface.co/datasets/cognitivecomputations/dolphin have benefited from Explanation Tuning, showing substantial improvements over traditional instruction-tuned models.
@@
[…]
@@ 5.1.3. Model Performance
Table 9. Comparative Performance of Various Models on Common Benchmarks: data from MobiLlama (Thawakar et al., 2024), OLMo (Groeneveld et al., 2024), and Llama 3.2.
@@
Accuracy is used as the evaluation metric in the table. Open Language Model (OLMo) (Groeneveld et al., 2024) is publicly available with its training data and code  [^9] https://allenai.org/olmo. MobiLlama (Thawakar et al., 2024) is a general-purpose SLM designed from scratch, available in 0.5B and 0.8B versions. It adopts a unique approach by using a shared FFN across all transformer blocks, enhancing efficiency. MobiLlama also show high efficiency on diverse hardware (Table 10). From Table 9, we can conclude that: (1) MobiLlama 0.5B and 0.8B demonstrate that a shared FFN design can facilitate excellent performance in SLMs with fewer than 1B parameters, even rivaling some models in the 1B-3B range. (2) The performance of MobiLlama 1.2B and OLMo 1.2B illustrates that advanced SLM architectures incorporating high-quality data, SwiGLU, non-parametric layer normalization, RoPE, BPE tokenization, and a shared FFN design can achieve competitive results among models with 1B-3B parameters. (3) MobiLlama demonstrates that SLMs can significantly reduce resource consumption on low-end hardware devices, achieving comparable performance while using a smaller proportion of battery power, memory, and GPU utilization. (4) Popular techniques such as pruning, quantization, distillation, SFT, and DPO, utilized in Llama 3.2, have substantially enhanced SLM performance.
Insights: We draw several key insights from the development of generic-specific SLMs: Typical SLM architectures generally incorporate features such as GQA, gated FFN with SiLU activations, RMS normalization, deep and thin architectures, embedding sharing, layer sharing, and shared FFNs. Although these components are widely used, current research has not yet thoroughly explored their specific contributions within SLMs. The importance of data quality in SLM research is increasingly emphasized, often considered more critical than the quantity of data and model architectural configurations. For example, the Dolma project enhances data quality through multi-stage data filtering processes. Post-pretraining, meticulous fine-tuning is often required to enhance the safety of SLMs, involving strategies to better distill capabilities from LLMs. Common strategies include explanatory tuning, progressive learning, and prompt erasing.
@@
MentaLLaMA (Yang et al., 2024f) 7B; 13B 2023.9 Instruction Tuning (LLaMA2 (Touvron et al., 2023b)) Healthcare IMHI dataset RLHF; PEFT
@@
Rho-1 (Lin et al., 2024a) 1B, 7B 2024.4 TinyLlama-1.1B (Zhang et al., 2024i), Mistral-7B (Jiang et al., 2023) Science (Mathematics) OpenWebMath (Paster et al., 2024) Continual pretraining
@@
OceanGPT (Bi et al., 2023) 2B, 7B, 14B 2023.10 LLaMA2 (Touvron et al., 2023b) Science (Ocean) Open-access literature, DoINSTRUCT Continual pre-training, Instruction tuning
@@
MindLLM (Yang et al., 2023a) 1.3B, 3B 2023.10 From-scratch and Supervised Fine-tuning Law, Finance Pile (Gao et al., 2020), Wudao (Yuan et al., 2021), CBooks Train on Bilingual Mixture Data, SFT
@@ 5.2. Domain-Specific SLMs
The capability of LLMs to generate human-like text has significantly captured public interest, highlighting their potential in the field of general artificial intelligence. However, inefficiencies persist when integrating LLMs into specialized applications due to resource constraints. Unlike the need for extensive general knowledge and capabilities, domain-specific SLMs should focus on well-defined tasks and expertise pertinent to specific fields. For instance, specialized models can significantly impact biomedical research and healthcare by fine-tuning LLMs for interpretable mental health analysis, or assisting humans in legal dialogues and financial tasks through instruction tuning, showcasing their potential transformative influence. Given the limited number of SLMs specialized in specific domains, we demonstrate some existing SLMs individually across healthcare, science, finance, and law domains.

### v2

@@ 5.1. Generic-domain SLMs
SLMs, with fewer parameters than LLMs, enhance computational efficiency in pre-training, fine-tuning, and inference, reducing memory and energy demands—crucial for resource-limited environments. Their compact, localized nature boosts privacy, personalization, and response times, making them ideal for low-power edge devices. Therefore, SLMs are attracting increasing attention, and various models are being developed. Table 7 summarizes current representative generic-domain $42$ SLMs/SLM families. Although all chosen SLMs have similar architectures, they vary in specific training datasets and techniques, with some datasets not being openly available. Taking the latest Llama 3.2 1B models (AI, 2024) in Figure 19 as an example, its parameter size and use of filtered high-quality training data, pruning-based initialization, knowledge distillation pre-training tasks, and training techniques such as Supervised Fine-Tuning (SFT), Rejection Sampling (RS), and Direct Preference Optimization (DPO) distinguish it from others.
@@
5.1.1. Architecture Design
From Table 7, we observe several trends in component choices for SLMs:
  1. (1)
    Recent SLMs frequently employ Grouped Query Attention (GQA) in self-attention mechanisms because it can reduce computational complexity. GQA achieves this by sharing query representations across multiple heads while keeping key and value representations separate. This approach aligns with the goals of SLM to enhance efficiency without compromising functionality.
  2. (2)
    The choice of activation function should balance model capability and efficiency. ReLU, known for its efficiency, introduces greater sparsity to the model, which facilitates faster coefficient calculations for inference acceleration. In contrast, SwiGLU’s parameters are learned during training, allowing the model to dynamically adapt to diverse tasks and datasets, thereby enhancing model capabilities and establishing it as a state-of-the-art option. SiLU, situated between these two, is favored for its balance of computational efficiency and model performance.
  3. (3)
    RMS normalization is commonly used than layer normalization due to its reduced computational demands.
A basic introduction to these options is provided in Section 2. Apart from component choices, there are notable innovations in architecture for SLMs:
  • Mobilellm (Liu et al., 2024h) highlights that deeper models are more effective than wider ones for improving performance.
  • Embedding sharing (Zhang et al., 2022) is crucial as embedding layers often constitute over 20% of a model’s parameters—for example, with 512 dimensions and a 32k vocabulary, each layer holds 16M parameters in a 125M-parameter model. Smaller models often reuse these weights for both input and output layers, enhancing efficiency and compactness.
  • Layer sharing (Liu et al., 2024h) increases hidden layers in small Transformer models without additional storage costs.
@@ 5.1.1. Architecture Design
  • Architecture search ahead of pre-training. PhoneLM (Yi et al., 2024) proposes a principle for constructing on-device small language models: searching for a resource-efficient architecture on a given hardware to optimize the speed-capacity trade-off before pretraining. This approach inspires the tailored selection of architectural components for on-device SLMs, based on specific compositional requirements such as computing efficiency, model capability, and safety.
@@ 5.1.2. Training Datasets
From Table 7, we can observe a set of widely used training datasets in SLM development. We provide the details below:
  • Pile (Gao et al., 2020): It comprises 22 smaller, high-quality diverse corpora from various domains, such as Pile-CC, PubMed Central, ArXiv, GitHub, and FreeLaw, designed to offer a comprehensive foundation for language model training. The dataset contains 207 billion tokens and totals 825 GB.
@@
  • The Stack (Kocetkov et al., 2022): It contains 6 trillion tokens of source code from over 300 programming languages, useful for developing code-centric AI applications. Python-edu in smollm-corpus (Ben Allal et al., 2024) consists of Python files that are scored 4 or more by the educational code model and are extracted from the stack-v2-train dataset.
@@
  • FineWeb (Penedo et al., 2024), a large-scale (15-trillion tokens, 44 TB disk space) dataset for LLM pretraining. FineWeb is derived from 96 CommonCrawl snapshots. FineWeb-Edu is a subset of FineWeb constructed using scalable automated high-quality annotations for educational value.
From the analysis of these datasets, we can derive several critical insights regarding the development of SLMs: (i) Data quality is crucial for training effective SLMs, involving sophisticated filtering like removing duplicates or irrelevant content, often with another LLM’s help. For example, the TinyStories corpus (Eldan and Li, 2023) is tailored for simplicity, ideal for training models to handle straightforward narratives. RedPajama-V2 (Computer, 2023) uses the CCNet pipeline to process 30B documents, providing quality signals and IDs for creating a 20B deduplicated dataset. (ii) Code Data: Source code constitutes a significant component of valuable data for training models, particularly because of its structured nature and logical content. Training on code data enhances a model’s reasoning capabilities and supports its ability to generalize across multiple natural languages, which is crucial for applications requiring robust problem-solving and interpretation skills in diverse coding environments (MA et al., 2024; Aryabumi et al., 2024; Fu and Khot, 2022; Gunasekar et al., 2023)
@@ 5.1.3. Training Algorithms
  • Direct Preference Optimization (DPO) (Rafailov et al., 2024) presents a simpler alternative to RLHF for optimizing language models based on human preferences, preventing explicit reward modeling and reinforcement learning. Instead, DPO modifies log probabilities of responses with a dynamic weighting mechanism to prevent model degradation common in probability ratio-focused methods. The DPO loss function is:
@@
    $\mathcal{L}_{DPO}(\pi_{\theta};\pi_{\text{ref}})=-\mathbb{E}_{(x,y_{w},y_{l})\sim D}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)\right]$
@@
    where $\pi_{\theta}$ is the policy being optimized, $\pi_{\text{ref}}$ is the reference policy, $D$ includes tuples $(x,y_{w},y_{l})$, $\sigma$ is the sigmoid function, and $\beta$ scales the log-ratios between $\pi_{\theta}$ and $\pi_{\text{ref}}$, guiding the model towards human-preferred outputs.
@@
  • Reinforcement Learning from Contrast Distillation (RLCD) (Yang et al., 2024b) aims to calibrate generative SLMs/LLMs towards embodying harmless and beneficial characteristics. The process starts with an unaligned LM and initial prompts, which are modified into two variants $p+$ and $p-$, intended to promote and suppress, respectively, attributes like helpfulness and harmlessness. Upon inputting these prompts, the LM generates outputs $o+$ and $o-$, with $o+$ automatically designated as the preferred response. This automation speeds up training by avoiding additional evaluative scoring. The training continues under the RLHF framework.
  • Conditioned Reinforcement Learning Fine-Tuning (C-RLFT), by OpenChat (Wang et al., 2024d), enhances model performance by incorporating low-quality data during SFT. C-RLFT leverages varied data qualities with simple rewards (e.g., expert data at 1 credit, sub-optimal at 0.1), using distinct prompt tokens to condition data sources, eliminating costly human feedback. Similarly, Data Mix (Pfeiffer et al., 2024) trains on English text in three stages, reducing noisy web data progressively in each stage in favor of higher-quality data.
@@
  • Explanation Tuning, proposed by Orca (Mukherjee et al., 2023), addresses the limitations of standard instruction-based fine-tuning, which often restricts SLMs to style imitation rather than reasoning. It uses system prompts with instructions to direct GPT-4 to produce detailed explanations or perform step-by-step reasoning. The resulting instructions and the responses are used as a dataset for fine-tuning SLMs to have better ability of reasoning.
@@
  • Maximal Update Parameterization ($\mu$P) optimizes control initialization, layer-wise learning rates, and activation magnitudes to ensure stable training regardless of model layer widths. This method enhances training stability and allows the same optimizer settings, especially learning rates, to be used across different model scales. For instance, Cerebras-GPT (Soboleva et al., 2023) employs $\mu$P to train its models efficiently.
@@ 5.1.4. Model Performance
Table 8. Performance of Various SLMs on Common Benchmarks: data from MobiLlama (Thawakar et al., 2024), OLMo (Groeneveld et al., 2024), and Llama 3.2.
@@
Accuracy is used as the evaluation metric in the table. Open Language Model (OLMo) (Groeneveld et al., 2024) is publicly available with its training data and code  [^6] https://allenai.org/olmo. MobiLlama (Thawakar et al., 2024) is a general-purpose SLM designed from scratch, available in 0.5B and 0.8B versions. It adopts a unique approach by using a shared FFN across all transformer blocks, enhancing efficiency. MobiLlama also show high efficiency on diverse hardware (Table 9).
From Table 8 and 9, we can conclude that: (1) MobiLlama 0.5B and 0.8B demonstrate that a shared FFN design can facilitate excellent performance in SLMs with fewer than 1B parameters, even rivaling some models in the 1B-3B range. (2) The performance of MobiLlama 1.2B and OLMo 1.2B illustrates that advanced SLM architectures incorporating high-quality data, SwiGLU, non-parametric layer normalization, RoPE, BPE tokenization, and a shared FFN design can achieve competitive results among models with 1B-3B parameters. (3) MobiLlama demonstrates that SLMs can significantly reduce resource consumption on low-end hardware devices, achieving comparable performance while using a smaller proportion of battery power, memory, and GPU utilization. (4) Popular techniques such as pruning, quantization, distillation, SFT, and DPO, utilized in Llama 3.2, have substantially enhanced SLM performance.
Insights: We draw several key insights from the development of generic-domain SLMs: Typical SLM architectures generally incorporate features such as GQA, gated FFN with SiLU activations, RMS normalization, deep and thin architectures, embedding sharing, layer sharing, and shared FFNs. Although these components are widely used, current research has not yet thoroughly explored their specific contributions within SLMs. The importance of data quality in SLM research is increasingly emphasized, often considered more critical than the quantity of data and model architectural configurations. Post-pretraining, meticulous fine-tuning is often required to enhance the safety of SLMs, involving strategies to distill capabilities from LLMs better. Common strategies include explanatory tuning, progressive learning, and prompt erasing.
@@
MentaLLaMA (Yang et al., 2024i) 7B; 13B 2023.9 Instruction Tuning (LLaMA2 (Touvron et al., 2023b)) Healthcare IMHI dataset RLHF; PEFT
@@
Rho-1 (Lin et al., 2024a) 1B; 7B 2024.4 TinyLlama-1.1B (Zhang et al., 2024j), Mistral-7B (Jiang et al., 2023) Science (Mathematics) OpenWebMath (Paster et al., 2024) Continual pretraining
@@
OceanGPT (Bi et al., 2023) 2B; 7B; 14B 2023.10 LLaMA2 (Touvron et al., 2023b) Science (Ocean) Open-access literature, DoINSTRUCT Continual pre-training, Instruction tuning
@@
MindLLM (Yang et al., 2023a) 1.3B; 3B 2023.10 From-scratch and Supervised Fine-tuning Law, Finance Pile (Gao et al., 2020), Wudao (Yuan et al., 2021), CBooks Train on Bilingual Mixture Data, SFT
@@ 5.2. Domain-Specific SLMs
The capability of LLMs to generate human-like text has significantly captured public interest, highlighting their potential in the field of general artificial intelligence. However, inefficiencies persist when integrating LLMs into specialized applications due to resource constraints. Unlike the need for extensive general knowledge and capabilities, domain-specific SLMs should focus on well-defined tasks and expertise pertinent to specific fields. For instance, specialized models can significantly impact biomedical research and healthcare by fine-tuning for interpretable mental health analysis, or assisting humans in legal dialogues and financial tasks through instruction tuning, showcasing their potential transformative influence. Given the limited number of SLMs specialized in specific domains, we demonstrate some existing SLMs individually across healthcare, science, finance, and law domains.